A man, Rajiv, and a woman, Mengqi, were both born deaf. Only Rajiv has now died of old age. Together, Rajiv and Mengqi had two sons (Aiden and Ramesh) who were also born deaf. Aiden is married to Prudence who has normal hearing but their son, Bolin, was born deaf. Prudence is currently pregnant with their second child. Ramesh is married to Eva who has normal hearing and is pregnant with their first child. In the broader population, the frequency of the particular recessive allele causing deafness in this family is 0.01.

a) Draw a human pedigree (including names) consistent with this family description and showing the inheritance of deafness. Use appropriate symbols to indicate affected individuals and those certain to be deafness-allele carriers. [8 marks]

b) Calculate the probability that the second child of Aiden and Prudence will be born deaf. Explain your answer. [3 marks]

c) Calculate the probability that the child of Ramesh and Eva will be born deaf. Explain your answer.
Answer #1

Answer a:)

Mangqi Rajiv Eva Ramesh Prudence Aiden Bolin

Answer b:)

In Aiden and Prudence cross:

Prudence genotypes = Aa+XX

Aiden genotype =aa+XY

Aa+XX                   x                              aa+XY

F1-

Gametes

a+X

a+Y

A+X

Aa+XX carrier

Aa+XY carrier

a+X

aa+XX deceased

aa+XY deceased

There are two carriers and two deceased in F1 generation in which one child has been produced and he is deceased. Therefore, amongst three children, only one child is being possible to have deceased.

Therefore, probability of that one child = 1/3 x 100

Therefore, probability of that one child = 33.3%

Answer c:)

In Ramesh and Eva cross:

Eva genotypes = AA/a +XX

Ramesh genotype = aa+XY

AA/a+XX                              x                              aa+XY

F1-

Gametes

a+X

a+Y

A+X

Aa+XX carrier 0%

Aa+XY carrier 0%

a+X or A+X

A/aa+XX deceased or carrier

50% possibility

A/aa+XY deceased or carrier

50% possibility

Two offspring who have 50% possibility of having disease and other two have 0% possibility of having disease.

Now overall possibility of having disease in offspring = (50+50+0+0) /4

Now overall possibility of having disease in offspring = 100/4

Now overall possibility of having disease in offspring = 25%

Therefore, 25% possibility of having a deceased child of Ramesh and Eva

"""""""""Please rate the answer""""""""""